x + 3y = 13
2x – 3y = –19
(–2, 5)
(–2, 4)
(4, –2)
(5, –2)

Answer by LinnW(1048)

x + 3y = 13
2x – 3y = –19
add the two equations.
3x = -6
divide each side by 3
x = -2
substitute -2 for x in x + 3y = 13
-2 + 3y = 13
add 2 to each side
3y = 15
divide by 3
y = 5
So x = -2 and y = 5 and the point is (-2,5)